OpenStudy (anonymous):

Dustin needed a 5-letter password to access his online banking account. He decided to use only the letters in his first name for the password. If each letter must be different, how many possible passwords can there be?

OpenStudy (anonymous):

A 6 × 5 × 4 × 3 × 2 B 6 × 5 × 6 × 5 × 6 C 6 × 6 × 6 × 6 × 6

OpenStudy (mathstudent55):

First, we need to assume that Dustin is Dustin's first name.

OpenStudy (mathstudent55):

The name Dustin has 6 letters. For the first letter of his password, he can choose one of 6 letters: \(6 ...\) For the second letter, he only has 5 left to choose from: \(6 \times 5...\) For the third letter of the password, he has 4 letters left to choose from: \(6 \times 5 \times 4...\) If you keep going this way, you get to \(6 \times 5 \times 4 \times 3 \times 2\)

OpenStudy (anonymous):

its a